OpenStudy (anonymous):

The following function defines a recursive sequence: f(0) = -4 f(1) = 12 f(n) = -3•f(n -1) - 2•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -4, 12, -28, 60, … -4, -12, -28, -60, … -4, 12, -18, 54, … -4, 12, -18, -54, …

OpenStudy (anonymous):

@IMStuck help?

OpenStudy (anonymous):

@tox you there ?

OpenStudy (imstuck):

I will take a look, ok?

OpenStudy (anonymous):

-4, 12, -28, 60

OpenStudy (anonymous):

can someone explain ?

OpenStudy (anonymous):

-12 * 3 - 2 * (-4) = -36 + 8 = -28 ; - 3 * (-28) - 2 (12) = 84 - 24 = 60

OpenStudy (anonymous):

got it thanks guys